# The Pivoting Framework: Frank-Wolfe Algorithms with Active Set Size Control

### v2

@@ 1.1 Contributions
Second, we establish a connection between PM and active set identification, the process of identifying the face containing a solution of (OPT) and not to be confused with the active set of FW algorithms. When the feasible region is a polytope and the minimizers lie in the relative interior of a face ${\mathcal{C}}^{*}$ of ${\mathcal{C}}$, Bomze et al., (2020) provided sufficient conditions that guarantee the existence of an iteration ${R}\in\{0,1,\ldots,T\}$, where $T$ is the number of iterations AFW is run for, s.t. for all $t\in\{{R},{R}+1,\ldots,T\}$, the active set ${\mathcal{S}}^{(t)}$ produced by AFW is contained in the optimal face ${\mathcal{C}}^{*}$, implying that ${\mathbf{x}}^{(t)}\in{\mathcal{C}}^{*}$. In the same setting, we prove that applying PM to AFW guarantees an active set size of at most $\dim({\mathcal{C}}^{*})+1$ for all iterations $t\in\{{R},{R}+1,\ldots,T\}$. [^1] The result by Bomze et al., (2020) is not limited to AFW. In Appendix C, we derive a similar result for BPFW. Since our result holds for a more general setting, we further improve the upper bound on the size of the active set of $n+1$ to $\dim({\mathcal{C}})+1$ for any compact and convex feasible region ${\mathcal{C}}$.
@@ 2 Preliminaries
For $n\in\mathbb{N}$, let $[n]:=\{1,\ldots,n\}$. Vectors are denoted in bold. Given ${\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{n}$, let ${\mathbf{x}}\geq{\mathbf{y}}$ denote that $x_{i}\geq y_{i}$ for all $i\in[n]$. Given ${\mathbf{x}}\in\mathbb{R}^{n}$, let $\tilde{{\mathbf{x}}}\in\mathbb{R}^{n+2}$ be defined as $({\mathbf{x}}^{\intercal},0,1)^{\intercal}\in\mathbb{R}^{n+2}$. We denote the $i$th unit vector of dimension $n$ by ${\mathbf{e}}_{i}\in\mathbb{R}^{n}$. Given a vector ${\mathbf{x}}\in\mathbb{R}^{n}$, denote its support by $\operatorname{supp}({\mathbf{x}}):=\{i\in[n]\mid x_{i}\neq 0\}$. Given a matrix $M\in\mathbb{R}^{m\times n}$, we refer to the $(i,j)$th entry of $M$ as $M_{i,j}$ and the $i$th row and column of $M$ as $M_{i,:}$ and $M_{:,i}$, respectively. Furthermore, let $M_{:i,:j}$ denote the restriction of matrix $M$ to rows $1,\ldots,i$ and columns $1,\ldots,j$. We define the matrix
@@ 3.1 Carathéodory-amenable algorithms
4       $(\boldsymbol{\alpha}^{(t+1)},{\mathcal{S}}^{(t+1)},{\mathbf{x}}^{(t+1)})\leftarrow{\textnormal{{CCU}}}(\boldsymbol{\alpha}^{(t)},{\mathcal{S}}^{(t)},{\mathbf{x}}^{(t)})$triangleright see Algorithm 3
@@
These updates are captured by the more general convex-combination updates (CCUs) formalized in Algorithm 3. We refer to algorithms that perform repeated CCUs as Carathéodory-amenable algorithms (CAs), see Algorithm 2. Several comments are warranted. First, the running examples of this paper, FW, AFW, and BPFW, are CAs. Second, despite the clear focus of this work on FW algorithms, the class of CAs is general enough to potentially capture other methods, even though we are currently unaware of any other applications. Finally, we highlight that a CCU does not require ${\mathbf{v}}$ to be obtained via an (exact) LMO. Thus, the lazified variants of FW, AFW, and BPFW are also CAs.
@@ 3.2 Convex-combination-agnostic properties
We now formalize the concept of algorithmic properties of CAs that remain unchanged when one convex representation is replaced by another. As demonstrated in Section 4, the pivoting meta-algorithm (PM) ensures the preservation of these properties.
@@
[…]
@@ 1  $M^{(0)}\leftarrow\left(\tilde{{\mathbf{x}}}^{(0)},D_{n}\right)\in\mathbb{R}^{(n+2)\times(n+2)}$
5       $(\boldsymbol{\beta}^{(t+1)},{\mathcal{T}}^{(t+1)},{\mathbf{x}}^{(t+1)})\leftarrow{\textnormal{{CCU}}}(\boldsymbol{\alpha}^{(t)},{\mathcal{S}}^{(t)},{\mathbf{x}}^{(t)})$triangleright see Algorithm 3
6       $(M^{(t+1)},\boldsymbol{\alpha}^{(t+1)},{\mathcal{S}}^{(t+1)})\leftarrow{\textnormal{{ASC}}}(M^{(t)},\beta^{(t+1)},{\mathcal{T}}^{(t+1)},{\mathcal{T}}^{(t+1)}\setminus{\mathcal{S}}^{(t)})$triangleright see Algorithm 5
@@
1  if ${\mathcal{D}}=\emptyset$ then
@@ 3       $Q\leftarrow N$
5       ${\mathbf{v}}\in{\mathcal{D}}$triangleright ${\mathbf{v}}$ is unique
@@ 10       $\theta^{*}\leftarrow-\mu_{k}/r_{k}\geq 0$
11       ${\boldsymbol{\lambda}}\in\Delta_{n+2}$ s.t. for all $i\in[n+2]$, $\lambda_{i}\leftarrow\begin{cases}\theta^{*},&\text{if}\ i=k\\ \mu_{i}+\theta^{*}r_{i},&\text{if}\ i\neq k\\ \end{cases}$
@@
    $N$ is invertible, $N_{n+1,:}\geq\mathbf{0}^{\intercal}$, and $N_{n+2,:}\geq\mathbf{1}^{\intercal}$.
@@ 5 Active set identification results
We will now present improvements to the bound $|{\mathcal{S}}^{(t)}|\leq n+1$ established for PM in Theorem 4.2. First, we refine the bound to $|{\mathcal{S}}^{(t)}|\leq\dim({\mathcal{C}})+1$ for any iteration $t$. Then, we establish that if there exists an iteration ${R}$ such that ${\mathcal{S}}^{(t)}\subseteq{\mathcal{C}}^{*}\subseteq{\mathcal{C}}$ for all $t\geq{R}$, then $|{\mathcal{S}}^{(t)}|\leq\dim({\mathcal{C}}^{*})+1$ instead of $|{\mathcal{S}}^{(t)}|\leq n+1$ for all $t\geq{R}$. This result is tied to so-called active set identification properties of some FW variants, known for a long time in specific settings (Guélat and Marcotte,, 1986), and recentlt generalized in Bomze et al., (2020) for AFW when ${\mathcal{C}}$ is a polytope and some other mild assumptions are met. [^2] The result of Bomze et al., (2020) is proved in their paper for AFW. In Appendix C, we prove that a similar result also holds for BPFW.
@@
Consider the optimization problem (OPT) with ${\mathcal{C}}\subseteq\mathbb{R}^{n}$ a polytope, $f\colon{\mathcal{C}}\to\mathbb{R}$ a convex and $L$-smooth function, and the set of optimal solutions $\operatorname{argmin}_{{\mathbf{x}}\in{\mathcal{C}}}f({\mathbf{x}})$ lying in the relative interior of a face ${\mathcal{C}}^{*}\subseteq{\mathcal{C}}$. This setting is particularly relevant to the FW community, encompassing applications such as sparse signal recovery, sparse regression, and sparse logistic regression, where FW variants construct iterates that are sparse convex combinations of vertices of the feasible region. Second, vanilla FW with line-search or short-step is known to exhibit zig-zagging behavior and, in some cases, converges at a rate of at most $\Omega(1/t^{1+\epsilon})$ for any $\epsilon>0$ (Wolfe,, 1970). This motivated the development of accelerated variants (Lacoste-Julien and Jaggi,, 2015; Garber and Meshi,, 2016; Braun et al.,, 2019; Combettes and Pokutta,, 2020; Garber and Wolf,, 2021) and the exploration of alternative step-sizes (Wirth et al., 2023b, ; Wirth et al., 2023c, ) to surpass the lower bound established by Wolfe, (1970). Finally, the identification of the optimal face ${\mathcal{C}}^{*}$ is a crucial step known as active set identification; not to be confused with the active sets maintained by the CA’s. Once ${\mathcal{C}}^{*}$ is identified, the optimization problem (OPT) simplifies to one over ${\mathcal{C}}^{*}$, which often has a much lower dimension compared to ${\mathcal{C}}$ (Bomze et al.,, 2019; 2020).
@@
We similarly establish the active set identification property for BPFW and its convex-combination agnosticity in Theorem C.4. We obtain the following active set identification result for PM applied to AFW and BPFW.
@@ 6 Numerical experiments
We will now provide a brief set of numerical experiments. We discuss numerical robustness and stability in Section D and more extensive experiments as well as a detailed description of the setup in Section E in the supplementary materials.
We assess our algorithm on efficiency in terms of function value and FW gap convergence, and sparsity of the obtained solutions, compared to the standard and lazified versions (see (Braun et al.,, 2017) for details) of AFW and BPFW, which also produce notably sparse solutions. Our algorithm is implemented in Julia (Bezanson et al.,, 2017) v1.9.2 and builds on the FrankWolfe.jl package (Besançon et al.,, 2022). The sparse linear systems are solved with the LU decomposition of the UMFPACK library (Davis,, 2004). Plots are log-linear, function values are shifted so that the minimum on each plot reaches $10^{-8}$. The prefix L- denotes the lazified version of an algorithm, the prefix P- for the PM variant.
@@ 6.2 Sparse signal recovery
We assess our algorithm on a sparse signal recovery problem:
@@ Acknowledgments
We thank Zev Woodstock for providing valuable feedback for an earlier version of this manuscript. Research reported in this paper was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID 390685689, BMS Stipend).
A Missing proofs
@@
The proof is organized as follows: We first prove Properties 46 depending on whether the if (Line 5) or the else (Line 5) clauses are executed. Then, we prove Properties 79. First, suppose that ${\mathcal{D}}=\emptyset$. Then:
@@
$\min$ $-{\mathbf{e}}_{n+3}^{\intercal}{\boldsymbol{\gamma}}$ (A.1)
@@
The result follows by combining Theorem 5.1 with Theorem 5.3 for AFW, and with Theorem C.4 for BPFW, and noting that ${\mathbf{x}}^{(t)}\in{\mathcal{C}}^{*}$ for ${\mathcal{C}}^{*}\in\operatorname{faces}({\mathcal{C}})$ implies ${\mathcal{S}}^{(t)}\subseteq{\mathcal{C}}^{*}$. ∎
B FW variants

### v3

@@ 1.1 Contributions
Second, we establish a connection between PM and active set identification, the process of identifying the face containing a solution of (OPT) and not to be confused with the active set of FW algorithms. When the feasible region is a polytope and the minimizers lie in the relative interior of a face $\mathcal{C}$$^{*}$ of $\mathcal{C}$}, Bomze et al., (2020) provided sufficient conditions that guarantee the existence of an iteration ${R}\in\{0,1,\ldots,T\}$, where $T$ is the number of iterations AFW is run for, s.t. for all $t\in\{{R},{R}+1,\ldots,T\}$, the active set ${\mathcal{S}}^{(t)}$ produced by AFW is contained in the optimal face $\mathcal{C}$$^{*}$, implying that ${\mathbf{x}}^{(t)}\in{\mathcal{C}}^{*}$. In the same setting, we prove that applying PM to AFW guarantees an active set size of at most $\dim({\mathcal{C}}^{*})+1$ for all iterations $t\in\{{R},{R}+1,\ldots,T\}$. [^1] The result by Bomze et al., (2020) is not limited to AFW. In Appendix 9, we derive a similar result for BPFW. Since our result holds for a more general setting, we further improve the upper bound on the size of the active set of $n+1$ to $\dim({\mathcal{C}})+1$ for any compact and convex feasible region $\mathcal{C}$}.
@@ 2 Preliminaries
For $n\in\mathbb{N}$, let $n]:=\{1,\ldots,n\}$. Vectors are denoted in bold. Given ${\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{n}$, let ${\mathbf{x}}\geq{\mathbf{y}}$ denote that $x_{i}\geq y_{i}$ for all $i\in[n]$. Given ${\mathbf{x}}\in\mathbb{R}^{n}$, let $\tilde{{\mathbf{x}}}\in\mathbb{R}^{n+2}$ be defined as $({\mathbf{x}}^{\intercal},0,1)^{\intercal}\in\mathbb{R}^{n+2}$. We denote the $i$th unit vector of dimension $n$ by ${\mathbf{e}}_{i}\in\mathbb{R}^{n}$. Given a vector ${\mathbf{x}}\in\mathbb{R}^{n}$, denote its support by $\operatorname{supp}({\mathbf{x}}):=\{i\in[n]\mid x_{i}\neq 0\}$. Given a matrix $M\in\mathbb{R}^{m\times n}$, we refer to the $(i,j)$th entry of $M$ as $M_{i,j}$ and the $i$th row and column of $M$ as $M_{i,:}$ and $M_{:,i}$, respectively. Furthermore, let $M_{:i,:j}$ denote the restriction of matrix $M$ to rows $1,\ldots,i$ and columns $1,\ldots,j$. We define the matrix
@@ 3.1 Carathéodory-amenable algorithms
4 $(\boldsymbol{\alpha}^{(t+1)},{\mathcal{S}}^{(t+1)},{\mathbf{x}}^{(t+1)})\leftarrow{\textnormal{{CCU}}}(\boldsymbol{\alpha}^{(t)},{\mathcal{S}}^{(t)},{\mathbf{x}}^{(t)})$ $\triangleright$ see Algorithm 3
@@
These updates are captured by the more general convex-combination updates (CCUs) formalized in Algorithm 3. We refer to algorithms that perform repeated CCUs as Carathéodory-amenable algorithms (CAs), see Algorithm 2. Several comments are warranted. First, the running examples of this paper, FW, AFW, and BPFW, are CAs. Second, despite the focus of this work on FW algorithms, the class of CAs is general enough to potentially capture other methods, such as simplicial decompositions (Bettiol et al.,, 2024) or the nearest-extreme point variant (Garber and Wolf,, 2021). We leave the applicability of our framework for algorithms forming iterates as linear or conic combinations (such as matching pursuit) to future research. Finally, we highlight that a CCU does not require $\mathbf{v}$} to be obtained via an (exact) LMO. Thus, the lazified variants of FW, AFW, and BPFW are also CAs.
@@ 3.2 Convex-combination-agnostic properties
We now formalize in the following definition the concept of properties of CAs that remain unchanged when one convex representation is replaced by another. As demonstrated in Section 4, the pivoting meta-algorithm (PM) ensures the preservation of these properties.
@@
${\textbf{4a)}}\ (\boldsymbol{\beta}^{(t+1)},{\mathcal{T}}^{(t+1)},{\mathbf{x}}^{(t+1)})\leftarrow{\textnormal{{CCU}}}(\boldsymbol{\alpha}^{(t)},{\mathcal{S}}^{(t)},{\mathbf{x}}^{(t)})$
$\begin{aligned} {\textbf{4b)}}\ (\boldsymbol{\alpha}^{(t+1)},{\mathcal{S}}^{(t+1)})\leftarrow&\ \text{constructed via any procedure}\\ &\text{that guarantees that}\\ &{\mathbf{x}}^{(t+1)}=\sum_{{\mathbf{s}}\in{\mathcal{S}}^{(t+1)}}\alpha_{\mathbf{s}}^{(t+1)}{\mathbf{s}},\\ &\text{where}\ \boldsymbol{\alpha}^{(t+1)}\in\Delta_{|{\mathcal{V}}|}\ \text{and}\\ &{\mathcal{S}}^{(t+1)}=\{{\mathbf{s}}\in{\mathcal{T}}^{(t)}\mid\alpha_{\mathbf{s}}^{(t+1)}>0\}.\end{aligned}$
@@ 1  $M^{(0)}\leftarrow\left(\tilde{{\mathbf{x}}}^{(0)},D_{n}\right)\in\mathbb{R}^{(n+2)\times(n+2)}$
5 $(\boldsymbol{\beta}^{(t+1)},{\mathcal{T}}^{(t+1)},{\mathbf{x}}^{(t+1)})\leftarrow{\textnormal{{CCU}}}(\boldsymbol{\alpha}^{(t)},{\mathcal{S}}^{(t)},{\mathbf{x}}^{(t)})$ $\triangleright$ see Algorithm 3
6 $(M^{(t+1)},\boldsymbol{\alpha}^{(t+1)},{\mathcal{S}}^{(t+1)})\leftarrow{\textnormal{{ASC}}}(M^{(t)},\beta^{(t+1)},{\mathcal{T}}^{(t+1)},{\mathcal{T}}^{(t+1)}\setminus{\mathcal{S}}^{(t)})$ $\triangleright$ see Algorithm 5
@@
1  if $\mathcal{D}$$\emptyset$ then
@@ 3 $Q\leftarrow N$
5 ${\mathbf{v}}\in{\mathcal{D}}$ $\triangleright$ $\mathbf{v}$} is unique
@@ 10 $\theta^{*}\leftarrow-\mu_{k}/r_{k}\geq 0$
[…]
@@
    NN is invertible, $N_{n+1,:}\geq\mathbf{0}^{\intercal}$, and $N_{n+2,:}\geq\mathbf{1}^{\intercal}$.
@@ 5 Active set identification results
We will now present improvements to the bound $|{\mathcal{S}}^{(t)}|\leq n+1$ established for PM in Theorem 4.2. First, we refine the bound to $|{\mathcal{S}}^{(t)}|\leq\dim({\mathcal{C}})+1$ for any iteration $t$. Then, we establish that if there exists an iteration R{R} such that ${\mathcal{S}}^{(t)}\subseteq{\mathcal{C}}^{*}\subseteq{\mathcal{C}}$ for all $t\geq{R}$, then $|{\mathcal{S}}^{(t)}|\leq\dim({\mathcal{C}}^{*})+1$ instead of $|{\mathcal{S}}^{(t)}|\leq n+1$ for all $t\geq{R}$. This result is tied to so-called active set identification properties of some FW variants, known for a long time in specific settings (Guélat and Marcotte,, 1986), and recently generalized in Bomze et al., (2020) for AFW when $\mathcal{C}$} is a polytope and some other mild assumptions are met. [^2] The result of Bomze et al., (2020) is proved in their paper for AFW. In Appendix 9, we prove that a similar result also holds for BPFW.
@@
Consider the optimization problem (OPT) with ${\mathcal{C}}\subseteq\mathbb{R}^{n}$ a polytope, $f\colon{\mathcal{C}}\to\mathbb{R}$ a convex and $L$-smooth function, and the set of optimal solutions $\operatorname{argmin}_{{\mathbf{x}}\in{\mathcal{C}}}f({\mathbf{x}})$ lying in the relative interior of a face ${\mathcal{C}}^{*}\subseteq{\mathcal{C}}$. This setting is particularly relevant to the FW community, encompassing applications such as sparse signal recovery, sparse regression, and sparse logistic regression, where FW variants construct iterates that are sparse convex combinations of vertices of the feasible region. Second, vanilla FW with line-search or short-step is known to exhibit zig-zagging behavior and, in some cases, converges at a rate of at most $\Omega(1/t^{1+\epsilon})$ for any $\epsilon>0$ (Wolfe,, 1970). This motivated the development of accelerated variants (Lacoste-Julien and Jaggi,, 2015; Garber and Meshi,, 2016; Braun et al.,, 2019; Combettes and Pokutta,, 2020; Garber and Wolf,, 2021) and the exploration of alternative step sizes (Wirth et al., 2023b, ; Wirth et al., 2023c, ) to surpass the lower bound established by Wolfe, (1970). Finally, the identification of the optimal face $\mathcal{C}$$^{*}$ is a crucial step known as active set identification; not to be confused with the active sets maintained by the CA’s. Once $\mathcal{C}$$^{*}$ is identified, the optimization problem (OPT) simplifies to one over $\mathcal{C}$$^{*}$, which often has a much lower dimension compared to $\mathcal{C}$} (Bomze et al.,, 2019, 2020).
@@
We similarly establish the active set identification property for BPFW and its convex-combination agnosticity in Theorem 9.4. We obtain the following active set identification result for PM applied to AFW and BPFW.
@@ 6 Numerical experiments
We will now provide a brief set of numerical experiments. We discuss numerical robustness and stability in Section 10 and more extensive experiments as well as a detailed description of the setup in Section 11 in the supplementary materials.
We assess PM on efficiency in terms of function value and FW gap convergence, and sparsity of the obtained solutions, compared to the standard and lazified versions (see (Braun et al.,, 2017) for details) of AFW and BPFW, which also produce notably sparse solutions. Our algorithm is implemented in Julia (Bezanson et al.,, 2017) v1.9.2 and builds on the FrankWolfe.jl package (Besançon et al.,, 2022). The sparse linear systems are solved with the LU decomposition of the UMFPACK library (Davis,, 2004). Plots are log-linear, function values are shifted so that the minimum on each plot reaches $10^{-8}$. The prefix L- denotes the lazified version of an algorithm, the prefix P- for the PM variant.
@@ 6.2 Sparse signal recovery
We assess PM on a sparse signal recovery problem:
@@ Acknowledgments
We thank Zev Woodstock for providing valuable feedback for an earlier version of this manuscript. Research reported in this paper was partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID 390685689, BMS Stipend). Mathieu Besançon was partially supported by MIAI at Université Grenoble Alpes (ANR-19-P3IA-0003).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes/No/Not Applicable] Yes, all relevant properties of our algorithms are discussed and presented rigorously. See Sections 210.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes/No/Not Applicable] Yes. See Sections 210.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes/No/Not Applicable] Yes.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes/No/Not Applicable] Yes, our theoretical results strive to be self-contained. See Sections 210.
    2. (b)
      Complete proofs of all theoretical results. [Yes/No/Not Applicable] Yes, see Appendix 7.
    3. (c)
      Clear explanations of any assumptions. [Yes/No/Not Applicable] Yes. See Sections 210.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes/No/Not Applicable] The instructions are in the README of the supplementary directory.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes/No/Not Applicable] Yes, everything available in the supplementary materials.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes/No/Not Applicable] Not Applicable
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes/No/Not Applicable] Yes, beginning of Section 11.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes/No/Not Applicable] Yes.
    2. (b)
      The license information of the assets, if applicable. [Yes/No/Not Applicable] Not applicable.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes/No/Not Applicable] Not applicable.
    4. (d)
      Information about consent from data providers/curators. [Yes/No/Not Applicable] Not applicable.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Yes/No/Not Applicable] Not applicable.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Yes/No/Not Applicable] Not applicable.
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Yes/No/Not Applicable] Not applicable.
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Yes/No/Not Applicable] Not applicable.
7 Missing proofs
@@
The proof is organized as follows: We first prove Properties 46 depending on whether the if (Line 5) or the else (Line 5) clauses are executed. Then, we prove Properties 79. First, suppose that $\mathcal{D}$$\emptyset$. Then:
@@
$\min\$$-{\mathbf{e}}_{n+3}^{\intercal}{\boldsymbol{\gamma}}$ (7.1)
@@
The result follows by combining Theorem 5.1 with Theorem 5.3 for AFW, and with Theorem 9.4 for BPFW, and noting that ${\mathbf{x}}^{(t)}\in{\mathcal{C}}^{*}$ for ${\mathcal{C}}^{*}\in\operatorname{faces}({\mathcal{C}})$ implies ${\mathcal{S}}^{(t)}\subseteq{\mathcal{C}}^{*}$. ∎
8 FW variants